OpenStudy (anonymous):

(3x tot he 2/3 power) ^-1

jimthompson5910 (jim_thompson5910):

\[\Large \left(3x^{\frac{2}{3}}\right)^{-1}\] \[\Large 3x^{-\frac{2}{3}}\] \[\Large \frac{3}{x^{\frac{2}{3}}}\]